# Revisiting the impact of neutrino mass hierarchies on neutrino mass constraints in light of recent DESI data

### v1

@@ III.2 Planck and DESI
We now combine Planck data with DESI BAO data. The constraints are summarized in Tab. 2 and the posterior corner plot is shown in Fig. 3. The inclusion of DESI BAO data tightens the Bayesian constraints for all hierarchies. The tightest constraints are obtained for the 1M approximation ($M_{\mathrm{tot}}<0.084\,\mathrm{eV}$) and the DM approximation ($M_{\mathrm{tot}}<0.086\,\mathrm{eV}$), which – unlike for Planck data – give very similar results despite the 1M hierarchy being only a crude approximation [36]. Note that our upper limits from Planck+DESI data are slightly higher than the ones in the DESI BAO baseline result, which assume the DM approximation ($M_{\mathrm{tot}}<0.082\,\mathrm{eV}$, [6]) since we do not include ACT lensing data. Next we adopt the physically motivated NH and IH, which impose effective lower priors as described above (horizontal dotted lines in Fig. 3). Similar to the case with Planck data alone, the posteriors get cut off at higher values of $M_{\mathrm{tot}}$ leading to looser constraints: $M_{\mathrm{tot}}<0.13\,\mathrm{eV}$ (NH) $M_{\mathrm{tot}}<0.16\,\mathrm{eV}$ (IH). In App. A, we explicitly explore the impact of this lower limit by imposing the “NH/IH prior” in an analysis assuming the DM approximation. We find that the DM approximation with NH/IH priors agrees well with the full NH/IH analyses for Planck+DESI data, reconfirming that the looser constraints from NH/IH are driven by the imposed lower limit.
@@
To gain a deeper understanding of the tight upper limits from Planck+DESI data, we compare the Bayesian constraints to frequentist constraints obtained with profile likelihoods (right column of Tab. 2 and Fig. 4). All assumed hierarchies give tight constraints, where the tightest is obtained with the crude 1M approximation ($M_{\mathrm{tot}}<0.05\,\mathrm{eV}$). Slightly looser constraints are obtained for IH ($M_{\mathrm{tot}}<0.06\,\mathrm{eV}$) and for DM and NH (both $M_{\mathrm{tot}}<0.07\,\mathrm{eV}$). The different neutrino mass orderings give rise to similar profile likelihoods, where the difference in the profile likelihoods and in the upper limits is similar to the level of the numerical noise in the minimizations. [^6] At this precision, the inferred upper limits depend sensitively on the $\Delta\chi^{2}$-value around $0.05-0.07\,\mathrm{eV}$ and the width of the extrapolated parabola, which introduces some uncertainty due to the proximity to the physical boundary in $M_{\mathrm{tot}}=0$.
To disentangle the effect of imposing a lower physical bound $M_{\mathrm{tot}}>0.06\,\mathrm{eV}$ (NH) and $M_{\mathrm{tot}}>0.1\,\mathrm{eV}$ (IH) from the difference in the physical modeling, we conduct the Feldman-Cousins construction with these boundaries. For all neutrino models, this yields identical constraints quoted in parenthesis in Tab. 2: $M_{\mathrm{tot}}<0.11\,\mathrm{eV}$ (“NH lower bound”) and $M_{\mathrm{tot}}<0.14\,\mathrm{eV}$ (“IH lower bound”). Hence, imposing the lower limit dominates over the physical modeling differences in the different neutrino mass orderings.
The goodness of fit quantified by the absolute $\chi^{2}$-values under Planck+DESI data (bottom panel of Fig. 4) makes it evident that all neutrino mass orderings fit this data similarly well, as the profiles agree for $0.20<M_{\mathrm{tot}}<$ 0.2. This reaffirms that the Bayesian constraints for NH and IH give looser constraints than the DM approximation due to the effective prior imposed by neutrino oscillations experiments.
@@ IV Conclusions
For Planck data, we find that the DM approximation gives tighter constraints than the physically motivated NH and IH (Fig 1). This loosening of the constraints can be attributed to the lower bound imposed by neutrino oscillation experiments for NH and IH (Eq. 1), confirming previous results. This is corroborated by the frequentist profile likelihood analysis, which shows that the DM approximation fits the data similarly well as NH and IH (Fig. 2). The approximation with one massive and two massless neutrinos (1M), however, presents a worse fit than the other neutrino mass orderings and gives looser Bayesian and frequentist constraints. As already pointed out in the literature [28], the 1M hierarchy is not a good approximation and should be avoided.
@@
Acknowledgements: We are grateful for stimulating discussions with Andrea Cozzumbo, Elisa Ferreira, Tanvi Karwal, Julien Lesgourgues and Kushal Lodha. LH was supported by a William H. Miller fellowship. MK was supported by NSF Grant No. 2412361, NASA Grant No. 80NSSC24K1226, and the John Templeton Foundation. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (arch.jhu.edu), which is supported by the National Science Foundation (NSF) grant number OAC1920103.
@@ Appendix A Impact of the lower limit of $M_{\mathrm{tot}}$ imposed by neutrino oscillation experiments
The lower limit inferred from neutrino oscillation data, which imposes a flat prior with lower limit $M_{\mathrm{tot}}>0.06$ (“NH prior”) and $M_{\mathrm{tot}}>0.1$ (“IH prior”), makes it difficult to directly compare Bayesian constraints from NH/IH to the DM approximation, where for the latter we adopted the common lower limit of the prior $M_{\mathrm{tot}}>0$. A straightforward comparison can be achieved by imposing the same lower limits as in NH/IH to the common DM approximation (as was also explored in [1]). We show the resulting posteriors under Planck+DESI data in Fig. 5. As expected, the constraints of the DM approximation with NH prior ($M_{\mathrm{tot}}<0.12\,\mathrm{eV}$) agree with the analysis adopting the full NH mass ordering and the DM approximation with IH prior ($M_{\mathrm{tot}}<0.16\,\mathrm{eV}$) agrees well with the full IH mass ordering.

### v3

@@ IV.2 Planck and DESI
We now combine Planck data with DESI BAO data. The constraints are summarized in Tab. 2 and the posterior corner plot is shown in Fig. 3. The inclusion of DESI BAO data tightens the Bayesian constraints for all hierarchies. The tightest constraints are obtained for the 1M approximation ($M_{\mathrm{tot}}<0.083\,\mathrm{eV}$) and the DM approximation ($M_{\mathrm{tot}}<0.086\,\mathrm{eV}$), which – unlike for Planck data – give very similar results. Note that our upper limits from Planck+DESI data are slightly higher than the ones in the DESI BAO baseline result, which assume the DM approximation ($M_{\mathrm{tot}}<0.082\,\mathrm{eV}$, [6]) since we do not include ACT lensing data. Next we adopt the physically motivated NH and IH, which impose effective lower priors as described above (horizontal dotted lines in Fig. 3). Similar to the case with Planck data alone, the posteriors get cut off at higher values of $M_{\mathrm{tot}}$ leading to looser constraints: $M_{\mathrm{tot}}<0.13\,\mathrm{eV}$ (NH) and $M_{\mathrm{tot}}<0.16\,\mathrm{eV}$ (IH). In App. A, we explicitly explore the impact of this lower limit by imposing the “NH/IH prior” in an analysis assuming the DM approximation. We find that the DM approximation with NH/IH priors agrees well with the full NH/IH analyses for Planck+DESI data, reconfirming that the looser constraints from NH/IH are driven by the imposed lower limit.
@@
To gain a deeper understanding of the tight upper limits from Planck+DESI data, we compare the Bayesian constraints to frequentist constraints obtained with profile likelihoods (right column of Tab. 2 and Fig. 4). All assumed hierarchies give tight constraints, where the tightest is obtained with the 1M approximation ($M_{\mathrm{tot}}<0.05\,\mathrm{eV}$). Slightly looser constraints are obtained for IH ($M_{\mathrm{tot}}<0.06\,\mathrm{eV}$) and for DM and NH (both $M_{\mathrm{tot}}<0.07\,\mathrm{eV}$). Note that the upper limits for the IH are below the minimum allowed mass of $0.1\,\mathrm{eV}$ in the IH; thus, we will consider physically motivated lower bounds below. The different neutrino mass orderings give rise to similar profile likelihoods, where the difference in the profile likelihoods and in the upper limits is similar to the level of the numerical noise in the minimizations. [^6] At this precision, the inferred upper limits depend sensitively on the $\Delta\chi^{2}$-value around $0.05-0.07\,\mathrm{eV}$ and the width of the extrapolated parabola, which introduces some uncertainty due to the proximity to the physical boundary in $M_{\mathrm{tot}}=0$.
As discussed in Sec. II, the inclusion of BAO data partially breaks the geometrical degeneracy in the CMB between $M_{\nu}$, $\Omega_{m}$ and $H_{0}$. This leaves less wiggle room in the cosmological parameters to adapt the fit to CMB data and can explain the similar quality of fit to Planck+DESI data for all neutrino mass orderings. The impact of the different orderings on the geometry and background can explain the small differences in the four scenarios.
To disentangle the effect of imposing a lower physical bound $M_{\mathrm{tot}}>0.06\,\mathrm{eV}$ (NH) and $M_{\mathrm{tot}}>0.1\,\mathrm{eV}$ (IH) from the difference in the physical modeling, we conduct the Feldman-Cousins construction with these boundaries. For all neutrino models, this yields identical constraints quoted in parentheses in Tab. 2: $M_{\mathrm{tot}}<0.11\,\mathrm{eV}$ (“NH lower bound”) and $M_{\mathrm{tot}}<0.14\,\mathrm{eV}$ (“IH lower bound”). Hence, imposing the lower limit dominates over the physical modeling differences in the different neutrino mass orderings.
The goodness of fit quantified by the absolute $\chi^{2}$-values under Planck+DESI data (bottom panel of Fig. 4) makes it evident that all neutrino mass orderings fit the data similarly well, as the profiles agree for $0<M_{\mathrm{tot}}<0.2\,\mathrm{eV}$, with the 1M approximation presenting a slightly worse fit. This reaffirms that the Bayesian constraints for NH and IH give looser constraints than the DM approximation due to the effective prior imposed by neutrino oscillations experiments.
@@ V Conclusions
For Planck data, we find that the DM approximation gives tighter constraints than the physically motivated NH and IH (Fig 1). This loosening of the constraints can be attributed to the lower bound imposed by neutrino oscillation experiments for NH and IH (Eq. 1), confirming previous results. This is corroborated by the frequentist profile likelihood analysis, which shows that the DM approximation fits the data similarly well as NH and IH (Fig. 2). The approximation with one massive and two massless neutrinos (1M), however, presents a worse fit than the other neutrino mass orderings and gives looser Bayesian and frequentist constraints. Despite [44, 30, 45, 46, 47] pointing out that the 1M approximation leads to significant changes in the background and perturbations compared to NH/IH/DM, the 1M approximation is often set as a default configuration in samplers like MontePython and Cobaya [89]. Our exploration confirms that the 1M approximation should be avoided and replaced by the DM approximation in parameter inference.
@@
Data availability: The data and code that support the findings of this article are openly available [90, 91, 92].
Acknowledgements: We are grateful for stimulating discussions with Graeme Addison, Andrea Cozzumbo, Elisa Ferreira, Wayne Hu, Tanvi Karwal, Julien Lesgourgues and Kushal Lodha. LH was supported by a William H. Miller fellowship. MK was supported by NSF Grant No. 2412361, NASA Grant No. 80NSSC24K1226, and the John Templeton Foundation. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (arch.jhu.edu), which is supported by the National Science Foundation (NSF) grant number OAC1920103.
@@ Appendix A Impact of the lower limit of $M_{\mathrm{tot}}$ imposed by neutrino oscillation experiments
The lower limit inferred from neutrino oscillation data, which imposes a flat prior with lower limit $M_{\mathrm{tot}}>0.06\,\mathrm{eV}$ (“NH prior”) and $M_{\mathrm{tot}}>0.1\,\mathrm{eV}$ (“IH prior”), makes it difficult to directly compare Bayesian constraints from NH/IH to the DM approximation, where for the latter we adopted the common lower limit of the prior $M_{\mathrm{tot}}>0$. A straightforward comparison can be achieved by imposing the same lower limits as in NH/IH to the common DM approximation (as was also explored in [1]). We show the resulting posteriors under Planck+DESI data in Fig. 5. As expected, the constraints of the DM approximation with NH prior ($M_{\mathrm{tot}}<0.12\,\mathrm{eV}$) agree with the analysis adopting the full NH mass ordering and the DM approximation with IH prior ($M_{\mathrm{tot}}<0.16\,\mathrm{eV}$) agrees well with the full IH mass ordering.